Please assist with problem below and show steps. Thank you.
(16b - 2) -5(3b + 2) = -3

Answer by stanbon(75887)

(16b - 2) -5(3b + 2) = -3
------------------
16b - 2 - 15b - 10 = -3
-----
b - 12 = -3
b = 9
-----------
Cheers,
Stan H.
--------------